# Deep Learning-Enabled ISAC-OTFS Pre-equalization Design for Aerial-Terrestrial Networks

### v1

@@ Abstract
Orthogonal time frequency space (OTFS) modulation has been viewed as a promising technique for integrated sensing and communication (ISAC) systems and aerial-terrestrial networks, due to its delay-Doppler domain transmission property and strong Doppler-resistance capability. However, it also suffers from high processing complexity at the receiver. In this work, we propose a novel pre-equalization based ISAC-OTFS transmission framework, where the terrestrial base station (BS) executes pre-equalization based on its estimated channel state information (CSI). In particular, the mean square error of OTFS symbol demodulation and Cramér-Rao lower bound of sensing parameter estimation are derived, and their weighted sum is utilized as the metric for optimizing the pre-equalization matrix. To address the formulated problem while taking the time-varying CSI into consideration, a deep learning enabled channel prediction-based pre-equalization framework is proposed, where a parameter-level channel prediction module is utilized to decouple OTFS channel parameters, and a low-dimensional prediction network is leveraged to correct outdated CSI. A CSI processing module is then used to initialize the input of the pre-equalization module. Finally, a residual-structured deep neural network is cascaded to execute pre-equalization. Simulation results show that under the proposed framework, the demodulation complexity at the receiver as well as the pilot overhead for channel estimation, are significantly reduced, while the symbol detection performance approaches those of conventional minimum mean square error equalization and perfect CSI.
Index Terms:
OTFS, channel prediction, integrated sensing and communications, pre-equalization.
@@ I Introduction
As the key technology extending coverage and promoting ubiquitous connectivity, the aerial-terrestrial network has garnered widespread attention in recent years [1, 2, 3]. The operation of the aerial-terrestrial network depends on exceptional wireless communication and sensing capabilities, which enable data transmission, trajectory tracking, and intrusion detection [4], etc. However, the limited spectrum resources make achieving great connections and sensing in the aerial-terrestrial network challenging, necessitating higher spectral efficiency techniques to meet this demand. The integrated sensing and communications (ISAC) technology is regarded as a promising candidate [5, 6, 7]. ISAC technology utilizes a unified waveform to simultaneously achieve communication and sensing functions. That is to say, the waveform transmitted by the transmitter can be used for deliver data symbols to the communication receiver, while also being used for sensing parameter estimation at the sensing receiver. This unified approach shares hardware resources, spectrum resources, and signal processing algorithms, leading to benefits such as improved spectrum and energy efficiency, and reduced hardware costs. It should be noted that the contradiction between the communication signal and the sensing signal poses challenges in the ISAC waveform design [8]. Therefore, the ISAC waveform needs to achieve a balance between communication and sensing performance.
@@
Note that the aerial-terrestrial integrated network can involve the high-mobility scenarios. In such cases, the wireless channels may transition from stable to highly dynamic, or from time-invariant channel to fast time-variant channel, degrading communication and sensing performance [9]. For example, the performance of ISAC systems using multi-carrier waveforms (e.g., orthogonal frequency division multiplexing) can be severely degraded by inter-carrier interference [10]. Hence, how to achieve efficient communication and precise sensing under high mobility has emerged as a critical challenge for aerial-terrestrial ISAC networks.
@@
Among the possible candidates, orthogonal time frequency space (OTFS) exhibits high reliability compared to other commonly used waveforms [11]. Specifically, in OTFS modulation, the communication channel parameters and sensing parameters are of the same form, making it more suitable as an ISAC waveform. Hence, it received attention from researchers in recent years [12, 13, 14]. For example, in [15], an OTFS-based ISAC system was proposed, allowing for highly accurate range-velocity profiles without requiring large bandwidth transmissions or long-duration frames. The authors in [16] addressed the challenge of precisely estimating of user state information alongside reliable data transmission in high-mobility environments. The authors in [17] presented a deep learning-based algorithm for target detection using ISAC-OTFS waveforms, where this approach eliminates the need for pre-computing any statistical parameters. Given that ISAC systems typically has to sense of environmental parameters to perform specific tasks, some studies are shifting their focus from mere target detection to parameter estimation. For example, in [18], the authors minimized the bit error rate (BER) of the OTFS-based ISAC system while ensuring accurate sensing parameter estimation. The authors in [19] proposed a generalized likelihood ratio test (GLRT) based multi-target detection and delay-Doppler-angle estimation algorithm for multiple-input-multiple-output OTFS radar sensing. Based on this, the trade-off between sensing SNR and communication rate was achieved by jointly optimizing the spatial and delay-Doppler (DD)-domain waveforms. Note that the aforementioned works all assumed channel state information (CSI) is perfect; however, in practical applications, due to the high mobility of the aerial node, the AP may only be able to obtain outdated CSI, which will definitely degrade the communication performance.
@@
To address the above issue, some works have been done to refine CSI for OTFS communication systems. Specifically, the authors in [20] used the sensing echoes to assist in beam alignment and beam tracking, where Kalman filtering was employed to predict the angle-domain information for updating beamforming strategies. A convolutional long-short term memory recurrent neural network (CLRNet) was then proposed in [21] by introducing convolutional neural network modules for spatial feature extraction and long short-term memory (LSTM) modules for capturing temporal dependencies. In [22], based on deep learning, the roadside unit utilized vehicle angular parameters obtained from sensing echoes to predict beam directions for the next time instant, thereby achieving precise beam pairing with minimal latency. These works focused on channel prediction in the spatial domain, while there have been little works studying methods to obtain the delay and Doppler shift information of OTFS. The authors in [23] reconstructed the topology between communication nodes using the estimated sensing parameters of communication targets. Based on kinematic formulas, the reconstructed topology was then adopted to assist in predicting the complete channel, including delay and Doppler shift, without requiring channel estimation. In addition to leveraging topology relationships, the work in [24] capitalized on past channel information across multiple time slots to forecast the future communication precoding matrix, thereby minimizing the frame error rate and achieving ultra-reliable low latency communications. Note that the literature discussed above [20, 21, 22, 23, 24] primarily concentrates on enhancing communication performance, but the sensing aspect was not addressed. Additionally, these works focused on the beamforming or precoding design and such algorithms may not be suitable for the aerial receivers (e.g., unmanned aerial vehicle (UAV)) with limited signal processing capabilities. This is because the complex OTFS channel estimation and equalization algorithms still need to be designed at the aerial receiver, which leads to increased signal processing delays at the aerial receiver. In sum, under the outdated CSI scenario, the investigation of the OTFS waveforms to balance the performance between communication and sensing while ensuring low processing complexity at the UAV side still needs to be addressed.
@@
In this work, we explore the ISAC transmission framework for an OTFS-based aerial-terrestrial ISAC network. By taking into account the impact of channel outdatedness caused by high mobility and the limitations of communication signal processing capabilities at the UAV, a channel prediction-based pre-equalization (CPBPE) framework for ISAC waveforms is proposed to achieve a trade-off between communication and sensing performance. The main contributions of this paper are as follows:
@@
  • For the aerial-terrestrial network with one terrestrial access point (AP), one UAV UE, and multiple scatters, an ISAC transmission framework based on OTFS waveform is developed, which includes a pre-equalization module to reduce the complexity at the UE side. Under this transmission framework, the mean square error (MSE) of OTFS data symbols and the Cramér-Rao lower bound (CRLB) of sensing parameter estimation are analyzed. Based on this, we formulate a weighted MSE and CRLB optimization problem under the transmission power constraint to balance the communication and sensing performance.
@@
  • Since the original optimization problem is difficult to solve directly, and it also faces the issue of outdated CSI, by leveraging deep learning, we come up with a CPBPE framework. Specifically, the CPBPE framework includes a parameter-level channel prediction module with a low-dimensional input space, which directly corrects outdated CSI by utilizing the decoupled OTFS channel parameters, a CSI processing module used for the post-processing of the parameter-level channel data, and a residual-structured deep neural network (DNN)-based module for the pre-equalization design that considers both communication and sensing performance.
@@
  • Simulation results show that the proposed framework effectively enhances communication and sensing performance compared to pre-equalization design using outdated CSI and can approximate the performance achieved with true CSI. This indicates the robustness of the proposed framework for pre-equalization in high-mobility scenarios. Furthermore, we examine the processing complexity at the UAV, which has been greatly reduced.
@@
The remainder of this paper is organized as follows. Section II presents the system model, including the pre-equalization framework for ISAC signals. Section III analyzes the performance metrics of the proposed transmission framework and then formulates the optimization problem, considering the weighted communication and sensing performance. The CPBPE framework, including the channel prediction module, CSI processing module, and pre-equalization design module, is elaborated in Section  IV. In Section V, numerical results are discussed to evaluate the communication and sensing performance of the proposed algorithms under different parameter settings. Finally, the paper is concluded in Section VI.
@@ II System Model
We consider an ISAC-enabled aerial-terrestrial network under the high-mobility communication scenario, which consists of one AP, one UAV UE, and several scatters, as shown in Fig. 1. Therein, the AP communicates with the UE while sensing it. In this work, the OTFS waveform is employed for communication and sensing. Similar to [25, 26, 27], both the AP and UEs are assumed to be equipped with a single antenna.
Figure 1: An illustration of the aerial-terrestrial ISAC system model.
@@ II-A ISAC Pre-equalization Framework
To reduce spectrum resource occupancy and signal processing complexity of UE, we come up with an OTFS-based ISAC pre-equalization framework. As the name suggests, pre-equalization is a scheme that mitigates the impact of the channel at the AP in advance, thereby avoiding the complex channel estimation and equalization at the UE. Specifically, the comparison between the block diagrams of our ISAC pre-equalization framework and conventional OTFS system is illustrated in Fig. 2. Different from the conventional scheme, the OTFS modulated signal $\mathbf{s}_{t}$ is firstly pre-equalized by the pre-equalization matrix $\mathbf{P}_{t}$ at AP and then transmitted to UE, which directly performs symbol detection without channel estimation and channel equalization. In the meantime, the AP senses the UE based on the received echo signals. This means a a unified waveform is required to simultaneously fulfill both communication and sensing functions. In this work, we primarily focus on pre-equalization design at AP to accommodate both communication and sensing. The following design principles are taken into account during pre-equalization design:
@@ II-A1 Trade-off Between Communication and Sensing Performance
There is a contradiction between the randomness of the communication signal and the determinism of the sensing signal. Therefore, pre-equalization design needs to achieve a trade-off between communication and sensing performance under a specific requirement, thus improving the spectrum resource utilization.
@@ II-A2 Outdated CSI and Sensing Parameters
Note that waveform design generally requires the instantaneous CSI and sensing parameters, but in practical applications, often only outdated CSI and sensing parameters are available. Therefore, it is essential to design an algorithm that accounts for outdated CSI and sensing parameters to minimize the impact of pre-equalization design inaccuracies on communication and sensing performance.
@@ II-B Sensing Model
Let $\mathbf{s}_{t}$ denote the original OTFS modulated symbol in the DD domain. Therein, the modulated information symbols in $\mathbf{s}_{t}$ are assumed to be independent and identically distributed (i.i.d.) with unit power $\sigma^{2}_{s}$. Then, multiply $\mathbf{s}_{t}$ with the pre-equalization matrix $\mathbf{P}_{t}$ mentioned in II-A, the vector form of ISAC signal symbol in DD domain $\mathbf{x}_{t}$ can be obtained, i.e.,
@@
where $P_{S}$ is the total number of sensing paths, ${{h}}_{t,p_{S}}$ is the channel attenuation at the $p_{S}$-th path. ${{l}}_{t,p_{S}}$ and ${{k}}_{t,p_{S}}$ are the delay tap and Doppler tap of the $p_{S}$-th path, respectively. Their relationship with the real delay ${\tau}_{t,p_{S}}=\frac{2R_{t,p_{S}}}{c}$ and Doppler shift ${\nu}_{t,p_{S}}=\frac{2f_{0}v_{t,p_{S}}}{c}$ can be expressed as ${l}_{t,p_{S}}={{\tau}_{t,p_{S}}}{M\Delta f},{k}_{t,p_{S}}={{\nu}_{t,p_{S}}}{NT}$, respectively, where $R_{t,p_{S}}$ is the distance between the AP and the target, $f_{0}$ is the carrier frequency, $v_{t,p_{S}}$ is the radial velocity of the target relative to the AP, $M$ is the number of delay grids, and $N$ is the number of Doppler shift grids, $10\leq{{l}_{t,p_{S}}}\leq M-$ 1, $10\leq{{\nu}_{t,p_{S}}}\leq N-$ 1, $\Delta f$ is the subcarrier spacing, and $T$ is the time slot duration. In (4), $\boldsymbol{\Delta}(p)=\operatorname{diag}\left\{\eta^{0},\eta^{1},\ldots,\eta^{MN-1-l_{t,p}},\eta^{-l_{t,p}}\ldots,\eta^{-1}\right\}$ characterizes as the Doppler influence, where $\eta\triangleq e^{\frac{j2\pi k_{t,p}}{MN}}$.
@@
where $L$ denotes the length of the cyclic prefix. The $(i,j)$-th element of $\boldsymbol{\Psi}^{\left({{l}}_{t,p_{S}}\right)}$ is
@@
By performing Fourier transformation and symplectic finite Fourier transformation to the TD-domain received sensing signal $\mathbf{r}_{S,t}$, we have the DD-domain received sensing signal $\mathbf{y}_{S,t}$ at the $t$-th time slot as
@@ III-A Performance Metrics
III-A1 CRLB Analysis for Estimation of Sensing Parameters
@@
According to [30], the CRLB can be derived as the inverse of the Fisher Information Matrix (FIM) $\boldsymbol{\mathcal{I}}$, i.e.,
@@
where $\xi_{i}$ and $\xi_{j}$ represent the $i$-th and $j$-th element of $\boldsymbol{\xi}(\mathbf{y}_{S,t})$, respectively, $\mathbf{\Sigma}=\frac{1}{\sigma^{2}_{A}}\mathbf{I}$ is the covariance matrix of $\mathbf{z}_{S,t}$, and $\mathbf{H}_{S,t}^{DD,{p_{S}}}$ represents the channel matrix of the $p_{S}$-th sensing path. For natation simplicity, $\mathbf{H}_{S,t}^{DD,{p_{S}}}$ will be denoted as $\mathbf{H}^{{p_{S}}}$ by omitting the superscript $DD$ and subscripts $S$ and $t$, i.e.,
@@
The derivation of $\mathbf{H}_{k^{{}^{\prime}},k}^{p_{S}}$ in (22) w.r.t $\mathfrak{Re}\{h_{p_{S}}\}$, $\mathfrak{Im}\{h_{p_{S}}\}$, $\tau_{p_{S}}$, and $\nu_{p_{S}}$ can be then calculated as (26), (27), (28), and (III-A1), where $\mathcal{L}_{\mathrm{ISI}}\triangleq\left[M-\left\lceil\frac{\tau_{p_{S}}M}{T}\right\rceil,M-1\right]$.
@@
where the step (a) leverages the i.i.d. property of AWGN, with its covariance matrix represented as $\sigma^{2}_{A}\mathbf{I}$. The derivation of step (b) makes use of the circular nature of matrix trace. The derivation of step (c) is based on the assumption that the modulated information symbols in $\mathbf{s}_{t}$ are i.i.d. with unit power $\sigma^{2}_{s}$.
$\frac{\partial\mathbf{H}_{k,k^{{}^{\prime}}}^{p_{S}}\left[{l,l^{{}^{\prime}}}\right]}{\partial\mathfrak{Re}\{h_{p_{S}}\}}=\frac{1+j\mathfrak{Im}\{h_{p_{S}}\}}{NM}\frac{1-e^{j2\pi\left(\nu_{p_{S}}NT-k+k^{\prime}\right)}}{1-e^{j2\pi\frac{\left(\nu_{p_{S}}NT-k+k^{\prime}\right)}{N}}}\frac{1-e^{j2\pi\left(l^{\prime}-l+\tau_{p_{S}}M\Delta f\right)}}{1-e^{j2\pi\frac{\left(l^{\prime}-l+\tau_{p_{S}}M\Delta f\right)}{M}}}e^{-j2\pi\nu_{p_{S}}(\frac{l^{\prime}}{M\Delta f})}\begin{cases}e^{-j2\pi\left(\frac{k^{\prime}}{N}+\nu_{p_{S}}T\right)},\!\!&\!\!\!\!l^{{}^{\prime}}\in\mathcal{L}_{\mathrm{ISI}};\\ 1,&\mathrm{else}.\end{cases}$ (26)
$\frac{\partial\mathbf{H}_{k,k^{{}^{\prime}}}^{p_{S}}\left[{l,l^{{}^{\prime}}}\right]}{\partial\mathfrak{Im}\{h_{p_{S}}\}}=\frac{j+\mathfrak{Re}\{h_{p_{S}}\}}{NM}\frac{1-e^{j2\pi\left(\nu_{p_{S}}NT-k+k^{\prime}\right)}}{1-e^{j2\pi\frac{\left(\nu_{p_{S}}NT-k+k^{\prime}\right)}{N}}}\frac{1-e^{j2\pi\left(l^{\prime}-l+\tau_{p_{S}}M\Delta f\right)}}{1-e^{j2\pi\frac{\left(l^{\prime}-l+\tau_{p_{S}}M\Delta f\right)}{M}}}e^{-j2\pi\nu_{p_{S}}(\frac{l^{\prime}}{M\Delta f})}\begin{cases}e^{-j2\pi\left(\frac{k^{\prime}}{N}+\nu_{p_{S}}T\right)},&l^{{}^{\prime}}\in\mathcal{L}_{\mathrm{ISI}};\\ 1,&\mathrm{else}.\end{cases}$ (27)
$\frac{\partial\mathbf{H}_{k,k^{{}^{\prime}}}^{p_{S}}\left[{l,l^{{}^{\prime}}}\right]}{\partial\tau_{p_{S}}}=h_{p_{S}}{\sum\limits_{n,m}\,}e^{j2\pi{({\nu_{p_{S}}NT-k+k^{{}^{\prime}}})}\frac{n}{N}}e^{j2\pi{({l-l^{{}^{\prime}}-\tau_{p_{S}}M\Delta f})}\frac{m}{M}}(j2\pi m\Delta f)\frac{e^{j2\pi\nu_{p_{S}}{(\frac{l^{{}^{\prime}}}{M\Delta f})}}}{NM}\begin{cases}e^{-j2\pi{({\frac{k^{{}^{\prime}}}{N}+\nu_{p_{S}}T})}},&l^{{}^{\prime}}\in\mathcal{L}_{\mathrm{ISI}};\\ 1,&\mathrm{else}.\end{cases}$ (28)
$\frac{\partial\mathbf{H}_{k,k^{{}^{\prime}}}^{p_{S}}\left[{l,l^{{}^{\prime}}}\right]}{\partial\nu_{p_{S}}}=\frac{j2\pi h_{p_{S}}}{NM}\sum_{m}e^{j2\pi(l-l^{{}^{\prime}}-\tau_{p_{S}}M\Delta f)\frac{m}{M}}e^{j2\pi\nu_{p_{S}}\left(\frac{l^{{}^{\prime}}}{M\Delta f}\right)}$
×$\times\begin{cases}e^{-j2\pi\left(\frac{k^{{}^{\prime}}}{N}+\nu_{p_{S}}T\right)}\left[\sum_{n}e^{j2\pi(\nu_{p_{S}}NT-k+k^{{}^{\prime}})\frac{n}{N}}\left(\frac{l^{{}^{\prime}}}{M\Delta f}-T\right)+\sum_{n}nTe^{j2\pi(\nu_{p_{S}}NT-k+k^{{}^{\prime}})\frac{n}{N}}\right],&l^{{}^{\prime}}\in\mathcal{L}_{\mathrm{ISI}};\\ \sum_{n}e^{j2\pi(\nu_{p_{S}}NT-k+k^{{}^{\prime}})\frac{n}{N}}\frac{l^{{}^{\prime}}}{M\Delta f}+\sum_{n}nTe^{j2\pi(\nu_{p_{S}}NT-k+k^{{}^{\prime}})\frac{n}{N}},&\mathrm{else}.\end{cases}$ (29)

### v2

# Pre-equalization Design for ISAC-OTFS Air-Ground Transmission: A Deep Learning Approach

@@ Abstract
Despite the strong Doppler resilience capability, orthogonal time-frequency space (OTFS) modulation suffers from high channel estimation and equalization complexity at the receiver, hindering its applicability in air-ground transmission. In this paper, we propose a pre-equalization–based integrated sensing and communications-OTFS downlink transmission framework in which the terrestrial access point executes pre-equalization using the predicted channel state information (CSI), so that the unmanned aerial vehicle can perform direct symbol detection without channel equalization. In particular, the mean square error of OTFS symbol demodulation and Cramér-Rao lower bound of sensing parameter estimation are considered, with their weighted sum utilized as the metric for optimizing the pre-equalization matrix. To address the time-varying CSI, we develop a deep learning based framework composed of channel prediction and pre-equalization. In particular, a parameter-level channel prediction module is utilized to decouple OTFS channel parameters, and a low-dimensional prediction network is leveraged to correct outdated CSI, which is then used to initialize the input of the pre-equalization module. Finally, a dual-branch residual-structured deep neural network is cascaded to execute pre-equalization. Simulation results show that the proposed channel prediction-based pre-equalization framework significantly reduces receiver complexity and pilot overhead while achieving symbol detection performance close to minimum mean square error equalization with perfect CSI under high mobility, as well as substantially improving sensing accuracy.
@@ I Introduction
Efficient development of the low-altitude economy relies on close collaboration of aerial and terrestrial infrastructures [1]. As a result, air-ground transmission has attracted significant research attention in recent years [2, 3]. The two most critical perspectives in air-ground networks are exceptional wireless communication and sensing capabilities, which enable data transmission, intrusion detection, and trajectory planning [1, 4, 5], etc. However, the limited spectrum resources hinder simultaneous high-throughput communication and high-accuracy sensing in the air-ground network, necessitating higher spectral efficiency techniques to meet this demand. The integrated sensing and communications (ISAC) technology is regarded as a promising candidate [6, 7, 8, 9]. ISAC technology utilizes a unified waveform to simultaneously achieve communication and sensing functions. This unified approach shares hardware resources, spectrum resources, and signal processing algorithms, leading to benefits such as improved spectrum and energy efficiency, and reduced hardware costs.
@@
It is worth noting that the air-ground integrated network can involve the high-mobility scenarios. In such cases, the wireless channels may transition from stable to highly dynamic, or from time-invariant channel to fast time-variant channel, degrading communication and sensing performance [10]. For example, the performance of ISAC systems using multi-carrier waveforms (e.g., orthogonal frequency division multiplexing) can be severely degraded by inter-carrier interference [11]. Hence, how to achieve efficient communication and precise sensing under high mobility has emerged as a critical challenge for air-ground ISAC networks.
@@
Among the possible candidates, orthogonal time frequency space (OTFS) exhibits high reliability for communication and sensing compared to other commonly used waveforms [12]. Besides, OTFS offers a unique advantage for ISAC applications since its communication channel parameters and sensing parameters share the same form. Hence, it has received extensive attention from researchers in recent years [13, 14, 15]. For example, in [16], an OTFS-based ISAC scheme was proposed, allowing for highly accurate range-velocity profiles without requiring large bandwidth transmissions or long-duration frames. The authors in [17] addressed the challenge of precise estimation of user state information alongside reliable data transmission in high-mobility environments. A deep learning based target detector was designed in [18], which takes the delay–Doppler correlation matrix as input and learns to recognize target signatures through training over multiple channel and noise scenarios, thereby achieving adaptive target detection without threshold design. Given that ISAC systems typically has to sense environmental parameters to perform specific tasks, some studies shifted their focus from mere target detection to parameter estimation. For example, in [19, 20], the authors minimized the bit error rate (BER) of the OTFS-based ISAC system while ensuring accurate sensing parameter estimation. The authors in [21] proposed a generalized likelihood ratio test (GLRT) based multi-target detection and delay-Doppler-angle estimation algorithm for multiple-input-multiple-output OTFS radar sensing. Through jointly optimizing the spatial-delay-Doppler domain waveform, a better trade-off between sensing SNR and communication rate was achieved.
@@
Note that the aforementioned works were all based on perfect channel state information (CSI); however, in practical applications, the high mobility of the aerial nodes can result in significantly outdated CSI, which will definitely degrade the communication performance. To address this issue, some studies have attempted to predict the spatial domain channel to dynamically adjust beamforming. Specifically, the authors in [22] used the sensing echoes to assist in beam alignment and beam tracking, where Kalman filtering was employed to predict the angle-domain information for updating beamforming strategies. A convolutional long-short term memory recurrent neural network (CLRNet) was then proposed in [23] by introducing convolutional neural network modules for spatial feature extraction and long short-term memory (LSTM) modules for capturing temporal dependencies. In [24], based on deep learning, the roadside unit utilized vehicle angular parameters obtained from sensing echoes to predict beam directions for the next time instant, thereby achieving precise beam pairing with minimal latency. These works focused on channel prediction in the spatial domain, while there have been little works studying methods to obtain the delay and Doppler shift information of OTFS.
For delay-Doppler (DD) domain CSI, a novel sensing-assisted channel estimation algorithm was proposed in [25], where the BS firstly senses the topology of propagation environment, and then reconstructs the communication channel using the sensing parameters. Specifically, the reconstructed topology was adopted to assist in predicting the channel parameters, including delay and Doppler shift, based on kinematic formulas, such that only the fading of each path has to be estimated via pilots. On the other hand, the authors in [26] capitalized on past channel information across multiple time slots to forecast the next communication precoding matrix, resulting in lower frame error rate and achieving ultra-reliable low-latency communications. Although the aforementioned works enabled reduced pilot overhead, the significant channel equalization processing complexity of the OTFS receiver has not been addressed. This oversight may render them unsuitable for aerial receivers with limited signal processing capabilities. Besides, the sensing perspective was also not considered in the aforementioned works.
@@
In this work, we explore the air-ground OTFS-based ISAC downlink transmission system. By jointly considering the outdated channel caused by high mobility and the limited signal processing capabilities at the aerial receiver, a channel prediction-based pre-equalization (CP-PE) framework is proposed to achieve a better trade-off between communication and sensing performance. The main contributions of this paper are concluded as follows:
@@
  • For the air-ground network with one terrestrial access point (AP), one unmanned aerial vehicle (UAV) UE, and multiple scatters, an ISAC downlink transmission framework based on OTFS waveform is developed, which includes a pre-equalization module to reduce the complexity at the UE receiver side. Under this transmission framework, the mean square error (MSE) of OTFS data symbols and the Cramér-Rao lower bound (CRLB) of sensing parameter estimation are analyzed. Based on this, we formulate a weighted MSE and CRLB optimization problem under the transmission power constraint to balance the communication and sensing performance.
@@
  • To address the formulated complicated optimization problem as well as outdated CSI, we propose a CP-PE framework by leveraging the nonlinear mapping capability of deep learning. Specifically, the CP-PE framework consists of three modules. The first is a parameter-level channel prediction module with a low-dimensional input space, which directly corrects outdated CSI by utilizing the decoupled OTFS channel parameters. The second is a CSI processing module that performs post-processing on the parameter-level channel data. The third is a dual-branch residual-structured deep neural network (DNN)-based pre-equalization module, which jointly considers both communication and sensing performance.
@@
  • Simulation results show that the proposed framework effectively enhances communication and sensing performance compared to pre-equalization design using outdated CSI and can approximate the performance achieved with true CSI. This indicates the robustness of the proposed framework for pre-equalization in high-mobility scenarios. Furthermore, the processing complexity at the UAV is reduced by more than 75% compared to the conventional scheme.
@@
The remainder of this paper is organized as follows. Section II presents the system model, including the pre-equalization framework for ISAC signals. Section III analyzes the performance metrics of the proposed transmission framework and then formulates the optimization problem, considering the weighted communication and sensing performance. The CP-PE framework, including the channel prediction module, CSI processing module, and pre-equalization design module, is elaborated in Section  IV. In Section V, numerical results are discussed to evaluate the communication and sensing performance of the proposed algorithms under different parameter settings. Finally, the paper is concluded in Section VI.
@@ II System Model
We consider an ISAC-enabled air-ground downlink transmission system under the high-mobility communication scenario, which consists of one AP, one UAV UE, and several scatters, as shown in Fig. 1. Therein, the AP simultaneously senses and communicates with the UE via the OTFS waveform. As this study concentrates on pre-equalization design in the DD domain instead of spatial-domain, both the AP and UE are assumed to employ a single antenna, adopting the same setting as the studies in [27, 28, 29].
@@ II-A ISAC Pre-equalization Framework
To reduce spectrum resource occupancy and signal processing complexity of UE, we propose a pre-equalization framework for ISAC-OTFS downlink transmission systems. In particular, pre-equalization is a scheme that mitigates the impact of the channel at the transmitter (AP) side, thereby avoiding the complex channel estimation and equalization at the UE. Specifically, the comparison between the block diagrams of the pre-equalization based ISAC framework and conventional OTFS system is illustrated in Fig. 2. Different from the conventional scheme, the OTFS modulated signal $\mathbf{s}_{t}$ is firstly pre-equalized by the pre-equalization matrix $\mathbf{P}_{t}$ at AP and then transmitted to UE, which directly performs symbol detection without channel estimation and channel equalization. This approach can significantly reduce the signal processing complexity at the receiver. For the sensing process, only the aerial UE echo is considered, as clutter can be effectively suppressed using conventional moving target indication and clutter mitigation techniques [30, 31]. This means a unified waveform is required to simultaneously fulfill both communication and sensing functions. In this work, we primarily focus on pre-equalization design at AP to accommodate both communication and sensing. The following design principles are taken into account during pre-equalization design:
@@ II-A1 Trade-off Between Communication and Sensing Performance
There is a contradiction between the communication and sensing performance for a specific signal. Therefore, pre-equalization design needs to achieve a trade-off between communication and sensing performance under a specific requirement, thus improving the spectrum resource utilization.
@@ II-A2 Outdated CSI and Sensing Parameters
Note that waveform design generally requires the instantaneous CSI and sensing parameters, while in most practical applications, only outdated CSI and sensing parameters are available. Therefore, it is essential to design an algorithm that accounts for outdated CSI and sensing parameters to minimize the impact of pre-equalization design inaccuracies on communication and sensing performance.
Figure 1: An illustration of the air-ground ISAC system model.
@@ II-B Sensing Model
Let $\mathbf{s}_{t}$ denote the original OTFS modulated symbol in the DD domain. Therein, the modulated symbols in $\mathbf{s}_{t}$ are assumed to be independent and identically distributed (i.i.d.) with unit power $\sigma^{2}_{s}$. Through multiplying $\mathbf{s}_{t}$ with the pre-equalization matrix $\mathbf{P}_{t}$ mentioned in II-A, the vector form of DD-domain ISAC symbols $\mathbf{x}_{t}$ can be obtained as
@@
[…]
@@
where $L$ denotes the length of the cyclic prefix (CP). The $(i,j)$-th element of $\boldsymbol{\Psi}^{\left({{l}}_{t,p_{S}}\right)}$ is
@@
By performing Fourier transformation and symplectic finite Fourier transformation to the TD-domain received sensing signal $\mathbf{r}_{S,t}$, the DD-domain received sensing signal $\mathbf{y}_{S,t}$ at the $t$-th time slot is expressed as
@@ III-A Performance Metrics
III-A1 CRLB Analysis for Sensing Parameter Estimation
@@
where $\xi_{i}$ and $\xi_{j}$ represent the $i$-th and $j$-th element of $\boldsymbol{\xi}(\mathbf{y}_{S,t})$, respectively, $\mathbf{\Sigma}=\frac{1}{\sigma^{2}_{A}}\mathbf{I}$ is the covariance matrix of $\mathbf{z}_{S,t}$, and $\mathbf{H}_{S,t}^{DD,{p_{S}}}$ represents the channel matrix of the $p_{S}$-th sensing path. The derivation of the partial derivatives with respect to the relevant parameters $\frac{\partial\mathbf{H}_{S,t}^{DD,{p_{S}}}}{\partial\xi_{i}}$ and $\frac{\partial\mathbf{H}_{S,t}^{DD,{p_{S}}}}{\partial\xi_{j}}$ can be referred to in [35] and is omitted here for brevity. For natation simplicity, $\mathbf{H}_{S,t}^{DD,{p_{S}}}$ will be denoted as $\mathbf{H}^{{p_{S}}}$ by omitting the superscript $DD$ and subscripts $S$ and $t$, i.e.,
@@
where the step (a) leverages the i.i.d. property of AWGN, with its covariance matrix represented as $\sigma^{2}_{A}\mathbf{I}$. The derivation of step (b) makes use of the circular nature of matrix trace. The derivation of step (c) is based on the assumption that the modulated symbols in $\mathbf{s}_{t}$ are i.i.d. with unit power $\sigma^{2}_{s}$.